Hands-On Calendar has found that the cost per card of producing x pocket calendar cards is given by

 C'(x) = -0.02x + 50, for x ≤ 800,

where c(x) is the cost, in cents, per card. Find the total cost, in dollars, of producing 550 cards.


1
Expert's answer
2021-05-07T14:25:12-0400

 Ans:-

"C'(x) = -0.02x + 50," "for x \u2264 800"

integrating the above equation

"\Rightarrow \int C'(x)=\int-0.02x+50"


"\Rightarrow C(x)=-0.01x^2+50x"


to find the total cost, in dollars, of producing 550 cards then put "x=550"


"C(550)=-0.01(550)^2+50\times550=24475" $